Does the domain of the function f defined by f(x) =√(3-x) /√(x-2) is R-{2}

1
Expert's answer
2021-11-15T20:24:47-0500

No it is not R-{2}

"domain \space of\space\frac{\sqrt{3-x}}{\sqrt{x-2}}:\space= \begin{bmatrix}
 solution: & 2 \lt\ x\leq 3 \\
 interval\space notation: &( 2,3\rbrack
\end{bmatrix}"


"domain \ set\ builder \ notation=2 \lt\ x\leq 3"

"domain=(2,3]"